You found a treasure map and it said: à ¢ € œ Start at the well, go 100m straight south, then 30m 40 degrees north of west, then 25m straight east and finally 45m 75 degrees south of east.à ¢ €  How far from the well and in what direction is the hidden treasure? Use graphical method. Verify your answer in component method.


1
Expert's answer
2021-10-18T11:04:38-0400

121 m 85° S of E.


"A_E=0,\\
A_S=100,\\
B_E=-30\cos40°=-22.98,\\
B_S=-30\sin40°=-19.28,\\
C_E=25,\\
C_S=0,\\
D_E=45\cos75°=11.65,\\
D_S=45\sin75°=43.47.\\\space\\
R_E=A_E+B_E+C_E+D_E=13.67,\\
R_S=124.2,\\\space\\
R=\sqrt{R_E^2+R_S^2}=122\text{ m},\\\space\\
\theta=\tan^{-1}\frac{R_S}{R_E}=83.72°."